OpenStudy (anonymous):

what are the first five geometric sequence an= (3)^n-1, starting with ^n=1

OpenStudy (anonymous):

\[a_n = 3^{n-1}\] \[a_1 = 3^{0}=1\] \[a_2= 3^{1}=3\] \[a_3= 3^{2}=9\] \[a_4 = 3^{3}=27\] \[a_5 = 3^{5}81\]

OpenStudy (anonymous):

or \[a_n = 3^{n}-1\] \[a_1 = 3^{1}-1=3-1=2\] \[a_2 = 3^{2}-1=9-1=8\] \[a_3 = 3^{3}-1=27-1=26\] \[a_4 = 3^{4}-1=81-1=80\] \[a_n5 = 3^{5}-1=243-1=\]

OpenStudy (anonymous):

last one is \[a_5 = 3^{5}-1=243-1=242\]